OpenStudy (anonymous):

Write a recursive formula for the sequence 15, 26, 48, 92, 180, ... Then find the next term.

OpenStudy (anonymous):

A) \(a_n= 2a_n – 1 – 4,~where~a_1 = 15; 356\) B) \(a_n= 2a_n – 4,~where~a_1 = 15; 356\) C) \(a_n= 4 + 11 · 2^{n–1},~where~a_1 = 15; 356\) D) \(a_n= 3a_n – 1 – 19,~where~a_1 = 15; 356\)

OpenStudy (anonymous):